The data show the number of viewers for television stars with certain salaries. Find the regression​ equation, letting salary be the independent​ (x) variable. Find the best predicted number of viewers for a television star with a salary of ​$10 million. Is the result close to the actual number of​ viewers, 10.1The data show the number of viewers for television stars with certain salaries. Find the regression equation, letting salary ​million? Use a significance level of 0.05.
Answer #1

SOLUTION:

From given data,

The data show the number of viewers for television stars with certain salaries. Find the regression​ equation, letting salary be the independent​ (x) variable. Find the best predicted number of viewers for a television star with a salary of ​$10 million. Is the result close to the actual number of​ viewers, 10.1​million? Use a significance level of 0.05.

observation Salary (x)

Viewer

(y)

(x-\bar{x})2 (y-\bar{y})2 (x-\bar{x}) * (y-\bar{y})
1 106 6.6 7267.5625 0.16 85.25*(-0.4) = -34.1
2 13 4.2 60.0625 7.84 -7.75*(-2.8) = 21.7
3 12 7.6 76.5625 0.36 -8.75*(0.6) = -5.25
4 7 5.4 189.0625 2.56 -13.75*(-1.6) = 22
5 11 10.4 95.0625 11.56 -9.75*(3.4) = -33.15
6 5 7.4 248.0625 0.16 -15.75*(0.4) = -6.3
7 6 7.9 217.5625 0.81

-14.75*(0.9)

= -13.275

8 6 6.5 217.5625 0.25 -14.75*(-0.5)=7.375
Sum \Sigmax = 166 \Sigmay = 56 \Sigma(x-\bar{x})2 = 8371.5 \Sigma(y-\bar{y})2 =23.7 \Sigma(x-\bar{x}) * (y-\bar{y}) = -41

\bar{x} = \Sigmax /n = 166/8 = 20.75

\bar{y} = \Sigmay /n = 56/8 = 7

From above

Sxx = \Sigma(x-\bar{x})2 = 8371.5

Sxy = \Sigma(x-\bar{x}) * (y-\bar{y}) = -41

Slope = 31 = Sxy / Sxx =  -41 / 8371.5 = -0.004

Intercept = \beta _{0} = \bar{y} - 31(\bar{x})

\beta _{0} = 7 - (-0.00489)(20.75 ) = 7.101

What is the regression equation?

\hat{y} = \beta _{0} + 31x

\hat{y} = 7.101 -0.004 x

What is the best predicted number of viewers for a television star with a salary of $10 million?

Where,

x = 10 then

\hat{y} = 7.101 -0.004 (10)

\hat{y} = 7.061

The result close to the actual number of viewers, 10.1 million?

Where,

x = 10.1 then

\hat{y} = 7.101 -0.004 (10.1)

\hat{y} = 7.0606

The result is very close to the actual number of viewers of 10.1 million.

Please thumbs-up / vote up this answer if it was helpful. In case of any problem, please comment below. I will surely help. Down-votes are permanent and not notified to us, so we can't help in that case.